OpenStudy (nevillesnetwork):

Medal & Fan (plz help me asap!) Donna's company reimburses her expenses on food, lodging, and conveyance during business trips. The company pays $50 a day for food and lodging and $0.60 for each mile traveled. Donna drove 200 miles and was reimbursed $1,620. Part A: Create an equation that will determine the number of days x on the trip. (3 points) Part B: Solve this equation justifying each step with an algebraic property of equality. (6 points) Part C: How many days did Donna spend on this trip?

OpenStudy (studygurl14):

A: "$50 a day" --> 50x "$0.60 for each mile traveled" and "200 miles" --> 0.6(200) "reimbursed $1,620" --> = $1620 50x + 0.6(200) = 1620 Solve for x

OpenStudy (nevillesnetwork):

Thank you. Thats all I really needed, thanks again!

OpenStudy (studygurl14):

For Part B, make sure you are justifying each step in solving the equation

OpenStudy (nevillesnetwork):

Ok thanks :) can you double check it once I do it?

OpenStudy (studygurl14):

Sure :)

OpenStudy (nevillesnetwork):

Ok, give me a moment.

OpenStudy (nevillesnetwork):

50x + 0.6(200) = 1620 50x + 120 = 1620 <multiply 0.6(200) 50x + 120 - 120 = 1620 -120 <subtract 120 on both sides of the equation 50x = 1500 \[\frac{ 50x }{ 50x } = \frac{ 1500 }{ 50 }\] ^divide 50x on both sides of the equation x=30

OpenStudy (studygurl14):

The steps are good, but your reasons are not. You need the actual properties.

OpenStudy (nevillesnetwork):

What do you mean "reasons" like I need to write the reason for applying each step?

OpenStudy (studygurl14):

And, you aren't dividing 50x...you are dividing 50

OpenStudy (nevillesnetwork):

Thats what I meant, I'll fix it.

OpenStudy (nevillesnetwork):

Ah I get it, I need to write which property I used in each step, right?

OpenStudy (studygurl14):



OpenStudy (nevillesnetwork):

Oh ok, thanks again! I owe you big time! :)